OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2. simplify (1/2A + 1/2B) What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (anonymous):

@gezimbasha, Do you know how to do this? Im confused and Dont know how to start

OpenStudy (anonymous):

For the first part you have (1/2 * 1/2 + 1/2 * 1/2) = (1/4+1/4) = 1/2 For the second part we have. There are 6*6=36 possible outcomes for two cubes. Now we want to know what is the probability on the first cube to show one of the numbers from set {2,3,4} and we know that it is 1/2. The same can be said for the second cube, so the final result is: 1/2*1/2 =1/4=25%

OpenStudy (anonymous):

Wow great job! thank you!! :)

OpenStudy (anonymous):

you're welcome! just let me know when you need help ;)